the second angle is a triangle is one and a half times the measure of the first. the third is 15 degree more than half the first angle. find the measure of each angles in the triangle
Answer by stanbon(75887)

the second angle is a triangle is one and a half times the measure of the first. the third is 15 degree more than half the first angle. find the measure of each angles in the triangle
------
Let 1st be "x"
Then 2nd is 1.5x
And 3rd (x/2)+15
------------------------
Equation:
x + 1.5x + (x/2) + 15 = 180
----
2x + 3x + x = 360-30
6x = 330
x = 55
1st = x = 55 degrees
2nd = 82.5 degrees
3rd = 42.5 degrees
=======================
Cheers,
Stan H.
=============================